# 3D-ADAM: A Dataset for 3D Anomaly Detection in Advanced Manufacturing

### v1

@@ 3.2 Ground Truth Annotations
We provide accurate, high-detail ground truth annotations for each instance of a part, machine element and defect in dataset. Part annotations are provided as a segmentation mask, machine element annotations are a bounding box with a class label and defect annotations are a segmentation mask with a class label. These are demonstrated for an example part in Figure 4.
@@
For all cameras, a semi-automated segmentation method, based on Cutie cutie2023 , was employed to create foreground masks isolating the part in the scene. These were checked by hand and corrected where necessary. Defect and Machine Element ground truth annotations required expert knowledge of the manufacturing defects present in a part, and the machine elements on which they occurred. For this reason, all defect segmentation masks and machine element bounding boxes were annotated by hand on the 2D image data of each scan using the set captured from the MechMind Nano, as this sensor provided the highest level of detail of defects. A single expert annotator performed the annotations for all images in the MechMind Nano set, with verification and feedback support provided by HAL Robotics experts. Homographic transform methods were employed to complete the ground truth annotations. These methods calculated the transformation matrix between part masks from the Mech-Mind Nano image to the equivalent from each of the remaining sensors, applying this transform to the defect masks and machine element bounding boxes to propagate the annotations from the MechMind Nano image set to the remaining systems with these being checked by hand with minor corrections performed where necessary.
@@
4 Benchmark and Evaluation Methodology
We evaluate our dataset against a suite of models in a supervised learning setting at the tasks of 3D and RGB+3D Anomaly Detection and Localisation. For this task, we divide our dataset into 3 sets, an anomaly-free training set, containing all of the defect-free scans captured for the dataset. The annotated defective scans are partitioned into a test and validation set in a 60:40 share. To evaluate how existing 3D Anomaly Detection and Localisation models perform on our dataset, as well as assess the challenge presented by the dataset, we carried out a benchmark on the current state-of-the-art models. For this, we include the four currently leading or recently leading available models at 3D and RGB+3D Anomaly Detection and Segmentation task on the MVTec3D-AD Dataset mvtec2022 paperswithcodergb3d paperswithcode3d . These are Transfusion transfusion2024 , CPMF cpmf2024 , 3DSR 3dsr2024 and Shape-Guided shapeGuided2023 .
Evaluation Methodology.
To provide a strong foundation for comparing the performance of the models included within the benchmark set on our dataset at the tasks of 3D Anomaly Detection and localisation, standard evaluation metrics for anomaly detection are employed. This maintains a consistent evaluation criteria across models, which allows accurate comparison of each model’s performance with respect to the MVTec 3D-AD Dataset, as well as comparison of the performance of each respective model on our dataset. These standard criteria are defined on the image-level anomaly detection performance and the pixel-level anomaly localisation performance. For image-level performance, the Area Under Receiver Operator Curve (AUROC) metric is utilised, while for the pixel-level anomaly localisation the Area Under Per Region Overlap (AUPRO) metric is used.
5 Industry Expert Survey
@@
To validate the accuracy and suitability of our annotation methodology for industrial applications, we conducted a survey with industry experts, who were asked to re-annotate a subset of our data. Five industry experts from three different manufacturing organisations, across the industrial research and development, digital process manufacturing and robotics sectors from both private and public sector organisations, volunteered to take part in this survey. Each was provided with a randomised set of defective scans representing 1% of the total dataset, with the intention that the task should be completable by a single expert in a few hours. A detailed set of instructions for the labelling task was provided, also included in our github repository, as well as "good" examples for the images assigned, providing as-designed, defect-free examples of each part instance alongside the defective case. Other than those images included in the instruction guide, participants were not provided with any examples of annotated images from the core dataset, in order to minimise the degree of influence on their approach to annotations, and provide as close to a blind evaluation of the annotation process as possible. The ground truth annotations produced by experts were then evaluated for similarity to the equivalent ground truth annotations from the core dataset using a similar methodology as Section 4. However, given all labelling is conducted based on 2D image data, in place of the AUROC and AUPRO methods employed for Image-Level and Pixel-Level comparison respectively which is appropriate in 3D contexts, we instead use the Intersection over Union (IoU) criteria for image-level comparison and the PRO (Per Region Overlap) criteria, which is the core of the AUPRO method and more appropriate for 2D contexts.
Figure 5: Comparison of Pixel-Level annotations between dataset ground truth and labeller annotations. Light Blue: True Positive; Red: False Negative; and Green: False Postive.
@@
The results of the labelling survey demonstrate strong correlation between the annotations proposed within our dataset and those produced by industry expert surveys. Mean IoU score across all surveyed experts was 0.6 with a standard deviation of 0.04, indicating good image-level agreement on overall defect annotations for a given sample, with scores amongst labellers being tightly clustered. A mean PRO score of 0.76 with a standard deviation of 0.058 indicates strong pixel-level agreement on the shape and location of individual defects, with results again exhibiting low variability between labellers. Figure 5 provides a visual example where we have taken an instance where the task sets of two labellers overlapped for this single example. Alignment between labellers and the ground truth is strong, with labellers demonstrating strong correlation of true positive pixels, and tending to favour false positives over false negatives, thereby indicating that industry experts scrutinised a slightly more rigorous threshold for considering a given region to be defective in comparison to that proposed in our dataset.
6 Discussion
@@
The results of the benchmarking evaluation and are included in the supplementary material. We find that across the set of models tested with our dataset, the results indicate that the 3D-ADAM dataset provides a significantly greater challenge in comparison to the performance of the same models on the MVTec3D-AD dataset. The 3D-ADAM dataset presents additional new challenges in 3D Anomaly Detection, Localisation and Classification which are not currently facilitated by available datasets. Such challenges include few-shot and zero-shot multi-modal anomaly detection and localisation, multi-modal anomaly classification, and machine element feature classification. Additionally, Unsupervised Anomaly Detection tasks can be performed with our dataset, though this is not explored here and presents an exciting opportunity for future work. No current datasets exist which are appropriate for unsupervised settings in 3D and RGB+3D Anomaly Detection.
Contribution.
The 3D-ADAM dataset is demonstrably more challenging than those preceding it. It proposes a uniquely comprehensive range of both defect types, covering a breadth of industrial defect types with instances presented across a range of annotated machine elements, which is a crucial factor for manufacturing applications and something no prior dataset has attempted. This dataset also uniquely proposes images and 3D scans of machine parts captured in-the-wild from multiple distinct 3D imaging sensors, further exemplifying its suitability for developing solutions suitable for real-world industrial deployment in advanced manufacturing use cases.
Broader Impact and Ethical Concerns.
Given the importance of identifying defects in manufacturing environments as a sector-wide challenge, the development of novel solutions for anomaly detection based on 3D image data is crucial for delivering a new generation of technologies for fully automating the detection, classification, localisation and ultimately correction of manufactured parts, in-process. The results of our benchmarking evaluation demonstrate that current leading models are not well equipped for this task due to the limitations of existing datasets for training, testing and validation. Our model provides a significant increase in the challenge presented, bringing it in line with realistic real-world use-cases that would be faced on deployment, thereby contributing to the broader ambition of developing mature 3D anomaly detection solutions in line with what is currently available in 2D anomaly detection. We anticipate no concerns around security, personal data protection, violation of intellectual property rights or any other major risks with this work. All part models included are obtained from open-source and freely licensed repositories, the additive manufacturing process was conducted using commercial-grade systems and materials and scanning was conducted in a closed cell without people present. The industry expert survey was conducted anonymously and no details pertaining to the experts involved or their individual contributions we retained once the results of the survey had been obtained.
Limitations.
@@
We acknowledge that opportunities for improvement exist beyond what is proposed by our dataset. By composing our dataset exclusively of PLA-based additively manufactured parts, we present only one type of texture for models to identify surface defects upon. In many high-value advanced manufacturing applications, subtle differences in the presentation of certain defects may occur due to differences in material properties. The degree this would impact applications in real-world use cases has not been explored and presents an interesting opportunity for future study. Additionally, while the range of surface defects accounted for in our dataset is the most comprehensive of any dataset for 3D Anomaly Detection to date, there remains an array of potential defects which can occur in advanced manufacturing use-cases, well understood by experts in the manufacturing and materials science fields, which are not provided in this dataset. This presents an opportunity for further augmentations to accommodate a broader range of industrial applications. Lastly, our dataset does not provide full 3D views of the parts included, as some datasets which precede it do, however we do not consider this a significant limitation, as we consider the availability of such complete scans to be an unrealistic standard to expect for real-world applications and instead anticipate that the multi-view, multi-camera in-the-wild approach proposed is significantly more appropriate for such instances.
7 Conclusion
@@
In this work, we present 3D-ADAM, a novel dataset for 3D anomaly detection, which aims to present a challenge closely aligned with the scenarios encountered in real-world industrial applications for Advanced Manufacturing. 3D-ADAM is the largest dataset for this task to date, hosting the broadest range of defects of any dataset for 3D anomaly detection. 3D-ADAM uniquely provides machine element annotations for each defect, which no previous dataset has proposed, and is unique in presenting this challenge through a multi-view, multi-camera, in-the-wild dataset. 3D-ADAM provides a significantly more challenging dataset than any prior dataset, and is unique in having its defect annotation methodology verified through a survey of industry experts.
Acknowledgements and Disclosure of Funding
This research was enabled in part through the manufacturing and laboratory resources, as well as the input of industry expertise provided by HAL Robotics. We also thank Jose Hernandez of the National Skills Academy, and Michael Canavan and Lewes Glen of the Industrial Design team at the National Manufacturing Institute Scotland, Andrew Hamilton of the Digital Process Manufacturing Centre, and Thibault Schwartz of HAL Robotics for their contributions in our survey of industry experts. Paul McHard has received funding from HAL Robotics and the Royal Commission for the Exhibition of 1851 via an Industrial Fellowship.

### v2

@@ III-B Ground Truth Annotations
We provide accurate, high-detail ground truth annotations for each instance of a part, machine element and defect in the dataset. Part annotations are provided as a segmentation mask, machine element annotations are a bounding box with a class label and defect annotations are a segmentation mask with a class label. These are demonstrated for an example part in Figure 4.
@@
For all cameras, a semi-automated segmentation method, based on Cutie [32], was employed to create foreground masks isolating the part in the scene. These were checked by hand and corrected where necessary. Defect and Machine Element ground truth annotations required expert knowledge of the manufacturing defects present in a part, and the machine elements on which they occurred. For this reason, all defect segmentation masks and machine element bounding boxes were annotated by hand on the 2D image data of each scan using the set captured from the MechMind Nano, as this sensor provided the highest level of detail of defects. A single expert annotator performed the annotations for all images in the MechMind Nano set, with verification and feedback support provided by domain experts. Homographic transform methods were employed to complete the ground truth annotations. These methods calculated the transformation matrix between part masks from the Mech-Mind Nano image to the equivalent from each of the remaining sensors, applying this transform to the defect masks and machine element bounding boxes to propagate the annotations from the MechMind Nano image set to the remaining systems with these being checked by hand with minor corrections performed where necessary.
IV Benchmark and Evaluation
IV-A Evaluation Methodology
We evaluate our dataset on a suite of models in an unsupervised learning setting at the tasks of RGB, 3D and RGB+3D Anomaly Detection and Segmentation. An unsupervised training regime was selected as it presented the current strongest performing suite of state of the art models in the literature. For these tasks, we prepare our dataset in a suitable unsupervised training protocol, dividing into 3 sets; an anomaly-free training set, containing all of the defect-free scans captured for the dataset, with the annotated defective scans partitioned into a test and validation set in a 60:40 share. Partitioning is performed on a part-instance level such that view-point leakage is prevented within the protocol.
Using the anomalib library [33], a set of three foundational and leading available models for 2D Anomaly Detection were evaluated, namely PatchCore [30], which we also evaluate in 3D protocol, UniNet [34] and DinoMaly [35]. To evaluate how existing 3D Anomaly Detection and Segmentation models perform on our dataset, as well as assess the challenge presented by the dataset, our benchmark includes the evaluation of three foundational models for 3D Anomaly Detection, as well as current state-of-the-art models for both 3D and RGB+3D Anomaly Detection and Segmentation. For the foundational models, we evaluate CFA [36], PaDiM [37] and PatchCore [30], all of which were trained and tested using the anomalib library [33]. Further, we include the three currently leading or recently leading available models at 3D and RGB+3D Anomaly Detection and Segmentation tasks on the MVTec3D-AD Dataset [11]. These are Transfusion [29], 3DSR [27] and GLFM [28].
IV-B Evaluation Metrics
To provide a strong foundation for benchmarking the performance of the models on our dataset at the tasks of RGB, 3D and RGB+3D Anomaly Detection and Segmentation, standard evaluation metrics for anomaly detection are employed. This maintains a consistent evaluation criteria across models, which allows accurate assessment of each model’s performance on our dataset. These standard criteria are defined on the image-level anomaly detection performance and the pixel-level anomaly localisation performance. For detection and localisation tasks, the image-level and pixel-level Area Under Receiver Operator Curve (AUROC) metrics are utilised respectively. For models in which segmentation tasks were available, pixel-level anomaly segmentation was evaluated using the standard Area Under Per Region Overlap (AUPRO) metric. Further, for those models evaluated using the anomalib library [33], performance metrics in the form of both image and pixel level F1 scores are also provided.
@@ IV-B Evaluation Metrics
IV-C Results
Table IV shows the results of the evaluation of 2D Anomaly Detection techniques on the 3D-ADAM dataset, noting strong performance at both image-level detection (Image AUROC) and pixel-level localisation (Pixel AUROC), consistent across models. Image F1 is considerably stronger in image-level detection tasks, and this is consistent across 2D techniques.
Table V shows the results of the evaluation of 3D and RGB+3D Anomaly Detection and Segmentation techniques on 3D-ADAM. It is notable here that foundational models perform well at detection (Image AUROC) and localisation (Pixel AUROC), with a similar trend of considerably stronger precision and recall in image-level detection (Image F1) as compared to pixel-level localisation (Pixel F1). Evaluation of the three leading models for 3D Anomaly Detection and Segmentation showed that models struggled significantly across available metrics as compared to their respective published results on the MVTec3D-AD Dataset, with the best performance being obtained by the GLFM model. Missing values, indicated by ”—-” in Tab. V are attributable to the fact that the foundational models deployed through the anomalib library [33] do not perform segmentation tasks, and therefore no AUPRO metric is calculated. Similarly, F1 scores are not computed by the three leading models; the authors of these respective models do not provide such metrics in their published evaluation, and we aim to take an unbiased approach for the assessment of these leading models. As such, F1 metrics for both Image and Pixel-level are respectively not computed. Figure 5 shows a visualisation of the anomaly predictions obtained from the 3D implementation of PatchCore [30], demonstrating that the model struggles significantly in accurately identifying defects as provided in the ground truth masks. Further examples of prediction results across the suite of models deployed in our benchmark can be found in the accompanying video.
Figure 5: Comparison of ground truth annotations with anomaly detection results achieved by PatchCore.
We hold that the findings of this evaluation demonstrate that the 3D-ADAM dataset provides significantly greater challenge to leading models in the field compared to other SOTA datasets. We attribute this difference to the in-the-wild nature of our dataset, including the variations to ambient lighting conditions included in the model, as well as the superior, industry-standard set of defect annotations we provide. This evidence further supports the challenge posed by our dataset and the opportunity to train models suitable for real-world deployment.
TABLE IV: Results for 2D Anomaly Detection and Localisation models on 3D-ADAM
Model Image AUROC Image F1 Pixel AUROC Pixel F1
PatchCore 0.962 0.959 0.865 0.069
UniNet 0.924 0.923 0.899 0.053
DinoMaly 0.834 0.903 0.959 0.200
Mean 0.907 0.929 0.908 0.107
TABLE V: Results for 3D and RGB+3D models Anomaly Detection, Localisation and Segmentation tasks on 3D-ADAM
Model Image AUROC Image F1 Pixel AUROC Pixel F1 AUPRO
CFA 0.794 0.939 0.870 0.058 —–
PaDiM 0.964 0.975 0.846 0.041 —–
PatchCore 0.976 0.977 0.842 0.061 —–
3DSR 0.485 —– 0.494 —– 0.184
TransFusion 0.633 —– 0.644 —– 0.310
GLFM 0.791 —– 0.820 —– 0.603
IV-D Industry Expert Survey
@@
To validate the accuracy and suitability of our annotation methodology for industrial applications, we conducted a survey with industry experts, who were asked to re-annotate a subset of our data. Five industry experts from three different manufacturing organisations, across the industrial research and development, digital process manufacturing and robotics sectors from both private and public sector organisations, volunteered to take part in this survey. Each was provided with a randomised set of defective scans representing 1% of the total dataset, with the intention that the task should be completable by a single expert in a few hours. A detailed set of instructions for the labelling task was provided, as well as ”good” examples for the images assigned, providing as-designed, defect-free examples of each part instance alongside the defective case. Participants were not provided with any examples of annotated images from the core dataset, in order to minimise the degree of influence on their approach to annotations, and provide as close to a blind evaluation of the annotation process as possible. The ground truth annotations produced by experts were then evaluated for similarity to the equivalent ground truth annotations from the core dataset using a similar methodology as Section IV. However, given all labelling is conducted based on 2D image data, in place of the AUROC and AUPRO methods employed for Image-Level and Pixel-Level comparison respectively which is appropriate in 3D contexts, we instead use the Intersection over Union (IoU) criteria for image-level comparison and the PRO (Per Region Overlap) criteria.
@@
The results of the labelling survey demonstrate strong correlation between the annotations within our dataset and those produced by industry expert surveys. Mean IoU score across all surveyed experts was 0.6 with a standard deviation of 0.04, indicating good image-level agreement on overall defect annotations for a given sample, with scores among labellers being tightly clustered. A mean PRO score of 0.76 with a standard deviation of 0.058 indicates strong pixel-level agreement on the shape and location of individual defects, with results again exhibiting low variability between labellers. The accompanying video provides a visual example where we have taken an instance where the task sets of two labellers overlapped for this single example. Alignment between labellers and the ground truth is strong, with labellers demonstrating strong correlation of true positive pixels, and tending to favour false positives over false negatives, thereby indicating that industry experts scrutinised a slightly more rigorous threshold for considering a given region to be defective in comparison to that proposed in our dataset.
V Discussion
@@
We find that across the set of models tested with our dataset, the results indicate that the 3D-ADAM dataset provides a significantly greater challenge in comparison to the performance of the same models on the MVTec3D-AD dataset. The 3D-ADAM dataset presents additional new challenges in 3D Anomaly Detection, Localisation and Classification which are not currently facilitated by available datasets. Such challenges include few-shot and zero-shot multi-modal anomaly detection and localisation, multi-modal anomaly classification, and machine element feature classification. Additionally, while our benchmark explores unsupervised 2D and 3D Anomaly Detection techniques exclusively, this is in order to both explore the performance of current leading models in literature, as well as maintain a consistent protocol across the benchmark. Moreover, with the 3D-ADAM dataset we make available protocols such that Supervised, Few-Shot and Zero-Shot Anomaly Detection techniques may all be utilised with our dataset, and presents an exciting opportunity for future work. This goes beyond what is provided by currently available leading datasets.
V-A Limitations
@@
We acknowledge that opportunities for improvement exist beyond what is proposed by our dataset. By composing our dataset exclusively of PLA-based additively manufactured parts, we present only one type of texture for models to identify surface defects upon. In many high-value advanced manufacturing applications, subtle differences in the presentation of certain defects may occur due to differences in material properties. However, as discussed, the literature consensus is that geometry is of far greater consequence to the occurrence of defects than material within the same additive manufacturing process. The degree this would impact applications in real-world use cases has not been explored and presents an interesting opportunity for future study.
While the range of surface defects accounted for in our dataset is the most comprehensive of any dataset for 3D Anomaly Detection to date, there remains an array of potential defects which can occur in advanced manufacturing use-cases, well understood by experts in the manufacturing and materials science fields, which are not provided in this dataset. This presents an opportunity for further augmentations to accommodate a broader range of industrial applications. Lastly, our dataset proides a multi-view scan of each part included, but does not provide full 3D views. We do not consider this a significant limitation, as we consider the availability of such complete scans to be unrealistic in real-world applications and instead anticipate that the multi-view, multi-camera in-the-wild approach proposed is significantly more appropriate for such contexts.
VI Conclusion
@@
In this work, we present 3D-ADAM, a novel dataset for 3D Anomaly Detection, which aims to present a challenge closely aligned with the scenarios encountered in real-world industrial applications for Additive Manufacturing. 3D-ADAM is the largest dataset for this task to date, hosting the broadest range of defects of any dataset for 3D Anomaly Detection. 3D-ADAM uniquely provides machine element annotations for each defect, which no previous dataset has proposed, and is unique in presenting this challenge through a multi-view, multi-camera, in-the-wild dataset. 3D-ADAM provides a significantly more challenging dataset than any prior dataset, and is unique in having its defect annotation methodology verified through a survey of industry experts.
Given the importance of identifying defects in manufacturing environments as a sector-wide challenge, the development of novel solutions for anomaly detection based on 3D image data is crucial for delivering a new generation of technologies for fully automating the detection, classification, localisation and ultimately correction of manufactured parts, in-process. The results of our benchmarking evaluation demonstrate that current leading models are not well equipped for this task due to the limitations of existing datasets for training, testing and validation. Our dataset provides a significant increase in the challenge presented, bringing it in line with realistic real-world use-cases that would be faced on deployment, thereby contributing to the broader ambition of developing mature 3D anomaly detection solutions in line with what is currently available in 2D anomaly detection.